### v2

# Seungwon Baek b,c    P. Ko d    Yuji Omura e    and Chaehyun Yu

Abstract
In this paper, we study two Higgs doublet models with gauged $U(1)_{H}$ symmetry instead of the usual softly broken $Z_{2}$ symmetry, motivated by the excesses around 96 GeV reported by the CMS collaboration in the searches for light resonances decaying to two photons and two $\tau$’s. In this model, one Higgs doublet field is charged under the $U(1)_{H}$ symmetry, thereby one can avoid tree-level flavor changing neutral currents. The extra $U(1)_{H}$ gauge symmetry requires extra chiral fermions, in order to satisfy the anomaly-free conditions. We analyze the signals of the light resonances, taking into account the contribution of the extra fermions, and discuss the consistency with the experimental results in this model.
1 Introduction
The Standard Model (SM) has been established as a theory describing particle physics. Almost all predictions of the SM are consistent with experimental results and the Higgs particle was finally discovered at the LHC ATLAS:2012yve ; CMS:2012qbp . It is certain that there are still large uncertainties in some observables, so that new physics may exist in the energy region that can be explored at the LHC. The new model, however, should not drastically modify the SM predictions. A lot of candidates for new physics have been proposed, motivated by problems in the SM. The mysteries, especially, concerned with the origin of the electroweak (EW) scale and the vacuum structure of our universe seem to suggest new particles around the EW scale. Thus, some extensions of the SM model that reveal new aspects of our universe may be confirmed near future.
Recently, the CMS collaboration has reported some excesses in the diphoton channel around 96 GeV mass region, that may suggest interesting possibilities of the vacuum structure. The CMS collaboration has surveyed resonances that decay to two photons, and reported deviations from the expected signals. Based on the data at $\sqrt{s}=$ 8 TeV and 13 TeV with the integrated luminosity of $19.7$ fb-1 and $35.9$ fb-1, respectively, the result shows a resonance at $95.3$ GeV with a local significance of $2.8\sigma$. This can be described by a signal strength, CMS:2018cyk
$\mu_{\gamma\gamma}^{\rm CMS,previous}=\frac{\sigma^{\rm exp}(gg\to\ s\to\gamma\gamma)}{\sigma^{\rm SM}(gg\to H^{\rm SM-like}\to\gamma\gamma)}=0.6\pm 0.2.$ (1)
Here $s$ denotes the 96 GeV scalar boson responsible for the resonance while the $H^{\rm SM-like}$ is the hypothetical Higgs boson with the mass for the resonance. The analysis with full data Run 2 data set has been already reported by the CMS collaboration. The result shows an excess with a local significance of $2.9\sigma$ at $95.4$ GeV. This signal strength for the resonance is CMS:HIG-20-002
$\mu_{\gamma\gamma}^{\rm CMS}=\frac{\sigma^{\rm exp}(gg\to\ s\to\gamma\gamma)}{\sigma^{\rm SM}(gg\to H^{\rm SM-like}\to\gamma\gamma)}=0.33^{+0.19}_{-0.12}.$ (2)
The ATLAS collaboration has also reported their analysis for the di-photon channel with the full Run 2 data set and found a mild excess with a local significance of $1.7\sigma$ ATLAStalk which corresponds to the signal strength of Biekotter:2023oen
$\mu_{\gamma\gamma}^{\rm ATLAS}=0.18^{+0.10}_{-0.10}.$ (3)
The combined signal strength of the CMS and ATLAS results without possible correlation is Biekotter:2023oen
$\mu_{\gamma\gamma}^{\rm exp}=0.24^{+0.09}_{-0.08}$ (4)
at the mass of $95.4$ GeV. The CMS collaboration has also reported another resonance which may be a candidate for an extra scalar boson in the di-$\tau$ channel. The local significance at 95 GeV is about $2.6\sigma$ and the corresponding signal strength is CMS:2022goy
$\mu_{\tau\tau}^{\rm CMS}=\frac{\sigma^{\rm exp}(gg\to\ s\to\tau\tau)}{\sigma^{\rm SM}(gg\to H^{\rm SM-like}\to\tau\tau)}=1.2\pm 0.5,$ (5)
while there has been no analysis at ATLAS for the corresponding region so far. In addition, the LEP collaboration has announced a resonance at a similar invariant mass of a $b\bar{b}$ pair. The resonance in the $e^{+}e^{-}\to Zs\to Zb\bar{b}$ channel was observed with a local significance of $2.3\sigma$. The excess can be interpreted as the signal strength of Cao:2016uwt ; Azatov:2012bz
$\mu_{b\bar{b}}^{\rm LEP}=\frac{\sigma^{\rm exp}(e^{+}e^{-}\to Z(s\to b\bar{b}))}{\sigma^{\rm SM}(e^{+}e^{-}\to Z(H^{\rm SM-like}\to b\bar{b}))}=0.117\pm 0.057,$ (6)
where the mass of the resonance is about 98 GeV.
If those deviations are originated from new resonances that reside around 96 GeV, one of the good candidates is a neutral scalar originated from scalar fields that contribute to the EW symmetry breaking Biekotter:2023oen ; Biekotter:2019kde ; Biekotter:2019mib ; Biekotter:2020cjs ; Biekotter:2021qbc ; Heinemeyer:2021msz ; Biekotter:2022jyr ; Biekotter:2023jld ; Bhattacharya:2023lmu ; Banik:2023ntf . The possibility that extra scalar fields may exist has been discussed in many works. Adding extra scalar fields is surely one simple way to extend the SM without disturbing the anomaly-free conditions. Scalar fields may play a role in breaking extra gauge symmetries. In fact, extra neutral scalars that are related to gauge symmetry breaking and reside around 96 GeV have been proposed motivated by the excesses, and the signals have been studied in many works Kundu:2019nqo ; Cao:2019ofo ; Abdelalim:2020xfk ; Belyaev:2023xnv ; Aguilar-Saavedra:2023tql ; Dutta:2023cig ; Ellwanger:2023zjc ; Cao:2023gkc ; Arcadi:2023smv ; Mulaudzi:2023khg ; Liu:2024cbr ; Cao:2024axg ; Kalinowski:2024uxe ; Ellwanger:2024txc ; Ellwanger:2024vvs ; Arhrib:2024wjj ; Benbrik:2024ptw ; Lian:2024smg ; Khanna:2024bah ; BrahimAit-Ouazghour:2024img ; Ashanujjaman:2023etj ; Banik:2023vxa ; Coloretti:2023yyq ; Dong:2024ipo ; Borah:2023hqw ; Ahriche:2023hho ; Ahriche:2023wkj ; Chen:2023bqr ; Dev:2023kzu ; Wang:2024bkg ; YaserAyazi:2024hpj ; Gao:2024qag ; Haisch:2017gql ; Fox:2017uwr ; Biekotter:2021ovi ; Azevedo:2023zkg .
One of the simple extensions of the SM with extra scalar fields contributing to the EW symmetry breaking is two Higgs doublet models (2HDMs). In generic 2HDMs, both of the Higgs doublets can contribute simultaneously to fermion masses when they develop nonzero vacuum expectation values (VEVs). Since the fermion mass matrices and Yukawa couplings are not simultaneously diagonalizable, Higgs mediated flavor-changing neutral currents (FCNCs) will be present even at tree levels. Then a large portion of the parameter space in the generic 2HDMs would have already been excluded by stringent constraints from flavor physics such as $M^{0}-\overline{M^{0}}$ mixings (with $M=K,B_{d},B_{s}$), $B\to X_{s}\gamma$, $B_{s}\rightarrow\mu^{+}\mu^{-}$, $l\rightarrow l^{\prime}\gamma,l\rightarrow 3l^{{}^{\prime}}$ etc., to name a few.
The Higgs-mediated FCNC problem in generic 2HDMs can be cured a la by Natural Flavor Conservation Criterion by Glashow and Weinberg Glashow:1976nt . A simple realization of their criterion is to introduce $Z_{2}$ symmetry under which two Higgs doublets has different $Z_{2}$ parity. Then the SM chiral fermions will also carry different $Z_{2}$ charges. There are four different charge assignments for which Yukawa couplings of the SM fermions are allowed in such a way that the fermions of the same electric charge get their masses from one and the same Higgs doublet, either $H_{1}$ or $H_{2}$, but not from both. Namely 2HDMs that do not predict tree level FCNCs are classified into four types. In the Type-I 2HDM, only one Higgs field is coupled to the SM fermions. In the Type-II 2HDM, one Higgs field ($H_{1}$) couples with right-handed down-type quarks and charged leptons, and the other Higgs field ($H_{2}$) couples with right-handed up-type quarks and neutrinos. It is possible to consider a model where the couplings are flipped: $H_{1}$ couples with right-handed down-type quarks and neutrinos, and the $H_{2}$ couples with right-handed up-type quarks and charged leptons. We call this model the Type-Y 2HDM. In the fourth 2HDM, $H_{1}$ couples with quarks and $H_{2}$ couples with leptons. Then this $Z_{2}$ symmetry is assumed to be softly broken by dim-2 operator, $m_{12}^{2}H_{1}^{\dagger}H_{2}+H.c.$, in order to increase the masses of the scalars originated from $H_{1}$ and $H_{2}$.
In this paper, we consider gauged $U(1)_{H}$ symmetry instead of softly broken $Z_{2}$ symmetry, by assigning a non-vanishing $U(1)_{H}$ charge to $H_{2}$ Ko:2012hd ; Ko:2013zsa ; Ko:2014uka ; Ko:2015fxa . In addition, extra complex singlet scalar field, $\Phi$, is also introduced in order to break $U(1)_{H}$ spontaneously. As discussed in Ref. Ko:2012hd , there are several possible setups in 2HDMs with gauged $U(1)_{H}$. In this work, we concentrate on the Type-II 2HDM and the Type-Y 2HDM motivated by the excesses around 96 GeV. In Sect. 2, we discuss the setups of our 2HDMs. In Sect. 3, we summarize experimental constraints relevant to our models. In Sect. 4, the analyses concerned with the excesses are shown. Section 5 is devoted to summary. In Appendix A, the constraints from the condition for the scalar potential to be bounded from below is introduced. In Appendix B, the relevant effective couplings are summarized.
2 The 2HDM with gauged $U(1)_{H}$ symmetry
We consider 2HDMs with extra $U(1)_{H}$ gauge symmetry, where one of the Higgs fields, $H_{2}$, is charged under the $U(1)_{H}$ gauge symmetry, while the other $H_{1}$ is not charged. Some SM fermions may be also charged under $U(1)_{H}$ so that phenomenologically viable Yukawa couplings for the SM fermions are allowed by the presumed gauge symmetry. Below, we discuss the detail based on Refs. Ko:2012hd ; Ko:2013zsa ; Ko:2014uka ; Ko:2015fxa .
Due to the charge assignment to the scalar doublet fields, both $H_{1}^{\dagger}H_{2}$ and $(H_{1}^{\dagger}H_{2})^{2}$ terms are forbidden in the model. It has been found that this setup for the scalar potential is not excluded by experiments so far Jung:2023ckm . However, because there is no new scale in the potential, the masses of the new scalar bosons in the model, in particular, the mass of the charged Higgs boson, are of at most EW scale of the VEV $v\sim 246$ GeV. We add a new singlet scalar field $\Phi$ to the model, which gives rise to the $H_{1}^{\dagger}H_{2}$ term after symmetry breaking. Then both the charged Higgs and extra neutral Higgs bosons can be heavier than the SM Higgs boson.
$H_{1}$ $H_{2}$ $\Phi$
$SU(3)_{c}$ 1 1 1
$SU(2)$ 2 2 1
$U(1)_{Y}$ $1/2$ $1/2$ 0
$U(1)_{H}$ 0 11 1-1
Table 1: The charge assignment to the scalar fields. Irreducible representations of $SU(3)_{c}$ and $SU(2)$ are denoted by their dimensions in boldface.
The charge assignments under $G_{SM}\times U(1)_{H}$ gauge symmetry to the scalar fields are shown in Table 1. Then the renormalizable parts of the scalar potential in our model is
$V(H_{k},\Phi)$ $=$ $m_{k}^{2}H_{k}^{\dagger}H_{k}+{\lambda_{k}}(H_{k}^{\dagger}H_{k})^{2}+\lambda_{3}(H_{1}^{\dagger}H_{1})(H_{2}^{\dagger}H_{2})+\lambda_{4}|H_{1}^{\dagger}H_{2}|^{2}$ (7)
$+\ m_{\Phi}^{2}|\Phi|^{2}+\lambda_{\Phi}|\Phi|^{4}+\tilde{\lambda}_{k}|\Phi|^{2}H_{k}^{\dagger}H_{k}-\left\{{\sqrt{2}\mu_{\Phi}H_{1}^{\dagger}H_{2}\Phi}+H.c.\right\},$
where $k=1,2$. We emphasize that, compared to the usual 2HDMs, two operators are missing in our model due to the $U(1)_{H}$ gauge symmetry: the soft $Z_{2}$ breaking dim-2 operator, $m_{12}^{2}H_{1}^{\dagger}H_{2}$, and the $\lambda_{5}(H_{1}^{\dagger}H_{2})^{2}+h.c.$ term. Still the $H_{1}^{\dagger}H_{2}+h.c.$ term can be realized from the $H_{1}^{\dagger}H_{2}\Phi+h.c$ terms after the singlet field $\Phi$ develops a non-vanishing VEV.
The scalar fields are expanded around nonzero VEVs with only $SU(3)_{C}\times U(1)_{\rm em}$ remaining unbroken:
$\langle H_{k}\rangle=(0,v_{k}/\sqrt{2})^{\intercal},\ \ \langle\Phi\rangle=v_{\Phi}/\sqrt{2},$ (8)
and
$H_{k}=\left(\begin{array}[]{c}\phi_{k}^{+}\\ \frac{v_{k}}{\sqrt{2}}+\frac{1}{\sqrt{2}}(h_{k}+i\chi_{k}^{0})\end{array}\right),\quad\Phi=\frac{1}{\sqrt{2}}(v_{\Phi}+h_{\Phi}+i\chi_{\Phi}),$ (9)
where $v_{1}=v\cos\beta$, $v_{2}=v\sin\beta$, and $v=\sqrt{v_{1}^{2}+v_{2}^{2}}=246$ GeV.
After the electroweak (EW) and $U(1)_{H}$ gauge symmetry breaking, three CP-even neutral scalar fields, $h_{i}\equiv(h_{1},h_{2},h_{\Phi})$, mix among themselves, and the $3\times 3$ mass matrix is diagonalized by three physical states, $S_{i}\equiv(\tilde{h},h,H)$. $h$ is identified as the Higgs boson with the mass of 125 GeV while $\tilde{h}$ is as the $96$ GeV scalar boson. The remaining $H$ is an additional scalar boson. The mixing between three scalar bosons is defined by
$S_{i}=R_{ij}h_{j}$}, (10)
where the rotation matrix $R_{ij}$ is given by
$R=\left(\begin{array}[]{ccc}1&0&0\\ 0&\cos\alpha_{3}&\sin\alpha_{3}\\ 0&-\sin\alpha_{3}&\cos\alpha_{3}\end{array}\right)\left(\begin{array}[]{ccc}\cos\alpha_{2}&0&\sin\alpha_{2}\\ 0&1&0\\ -\sin\alpha_{2}&0&\cos\alpha_{2}\end{array}\right)\left(\begin{array}[]{ccc}\cos\alpha_{1}&\sin\alpha_{1}&0\\ -\sin\alpha_{1}&\cos\alpha_{1}&0\\ 0&0&1\\ \end{array}\right)$ (11)
with three mixing angles, $\alpha_{i}$ $(i=1,2,3)$.
The CP-odd states, $\chi=(\chi_{\Phi},\chi_{1},\chi_{2})$, also mix among themselves and yield two Nambu-Goldstone bosons, $G_{1}^{0}$ and $G_{2}^{0}$, and one pseudoscalar boson, $A$. The mixing angles are defined by $G_{i}=(V_{A})_{ij}\chi_{j}$, where $G=(G_{1}^{0},G_{2}^{0},A)$ and
$V_{A}=\left(\begin{array}[]{ccc}0&\cos\beta&\sin\beta\\ \cos\delta&\sin\beta\sin\delta&-\cos\beta\sin\delta\\ \sin\delta&-\sin\beta\cos\delta&\cos\beta\cos\delta\end{array}\right)$ (12)
with
$\cos\delta=\frac{v_{\Phi}}{\sqrt{v_{\Phi}^{2}+(v\cos\beta\sin\beta)^{2}}}.$ (13)
Then the mass of the pseudoscalar boson is obtained by the mixing matrix as
$m_{A}^{2}=\mu_{\Phi}\left(\frac{v_{1}v_{\Phi}}{v_{2}}+\frac{v_{1}v_{2}}{v_{\Phi}}+\frac{v_{2}v_{\Phi}}{v_{1}}\right).$ (14)
The three neutral gauge bosons mix among them after EW and $U(1)_{H}$ symmetry breaking, and their mass matrix is given by
$M^{2}=\frac{v^{2}}{8}\left(\begin{array}[]{ccc}g^{2}&-gg_{Y}^{2}&-2gg_{X}s_{\beta}^{2}\\ -gg_{Y}&g_{Y}^{2}&2g_{Y}g_{X}s_{\beta}^{2}\\ -2gg_{X}s_{\beta}^{2}&2g_{Y}g_{X}s_{\beta}^{2}&4g_{X}^{2}\left(s_{\beta}^{2}+\frac{v_{\Phi}^{2}}{v^{2}}\right)\end{array}\right),$ (15)
where $g,g_{Y}$, and $g_{X}$ are gauge couplings for $U(1)_{Y}$, $SU(2)_{L}$, and $U(1)_{H}$, respectively, and $s_{\beta}=\sin\beta$. A massless state is identified as the photon and the remaining two states mix with each other. Then the two massive bosons, $Z$ and $Z^{{}^{\prime}}$, are expressed as
$\hat{Z}$ $=$ $Z\cos\theta+Z^{\prime}\sin\theta,$ (16)
$\hat{Z}^{\prime}$ $=$ $Z\sin\theta-Z^{\prime}\cos\theta,$ (17)
where the $\hat{Z}$ and $\hat{Z}^{\prime}$ are two states after identifying the photon state. The mixing angle is defined by
$\sin\theta=\frac{\lambda}{(1+\lambda^{2})^{1/2}},$ (18)
where
$\lambda$ $=$ $\frac{\bar{g}^{2}-g_{H}^{2}+\left[(\bar{g}^{2}-g_{H}^{2})^{2}+4\bar{g}^{2}\bar{\delta}^{2}\right]^{1/2}}{2\bar{g}\delta},$ (19)
$\bar{g}^{2}$ $=$ $g_{Y}^{2}+g^{2},$ (20)
$g_{H}^{2}$ $=$ $4g_{X}^{2}\left(s_{\beta}^{2}+\frac{v_{\Phi}^{2}}{v^{2}}\right),$ (21)
$\bar{\delta}$ $=$ $2g_{X}s_{\beta}^{2}.$ (22)
$Q^{i}_{L}$ $u^{i}_{R}$ $d^{i}_{R}$ $L_{L}^{i}$ $e^{i}_{R}$ $\nu^{i}_{R}$
$SU(3)_{c}$ 3 3 3 1 1 1
$SU(2)$ 2 1 1 2 1 1
$U(1)_{Y}$ $1/6$ $2/3$ 1/3-1/3 1/2-1/2 1-1 0
$U(1)_{H}$ 1/3-1/3 $2/3$ 1/3-1/3 0 0 11
Table 2: The charge assignment to the SM fermions in the Type-II 2HDM.
$Q^{i}_{L}$ $u^{i}_{R}$ $d^{i}_{R}$ $L_{L}^{i}$ $e^{i}_{R}$ $\nu^{i}_{R}$
$SU(3)_{c}$ 3 3 3 1 1 1
$SU(2)$ 2 1 1 2 1 1
$U(1)_{Y}$ $1/6$ $2/3$ 1/3-1/3 1/2-1/2 1-1 0
$U(1)_{H}$ 0 11 0 0 1-1 0
Table 3: The charge assignment to the SM fermions in the Type-Y 2HDM.
The fermion sector depends on the $U(1)_{H}$ charge assignments to the SM fermions. It turns out that the Type-I 2HDM for the fermions can be constructed by adding right-handed neutrinos without any gauge anomalies Ko:2012hd ; Ko:2013zsa ; Ko:2014uka , while models of other types require extra fermions as well as right-handed neutrinos in order to cancel gauge anomalies Ko:2012hd ; Ko:2015fxa . In this paper we focus on the Type-II and Type-Y models which are relevant to the 96 GeV scalar resonance signals Biekotter:2019kde . In Tables 2 and 3, we show the $U(1)_{H}$ charge assignments to the SM fermions in the Type-II and Type-Y 2HDMs, respectively.
Then we obtain the same Yukawa interactions as those in the usual 2HDM in both types. In the Type-II 2HDM, the Yukawa interaction is
${\cal L}_{\rm Yukawa}^{II}=-Y^{ij}_{u}\overline{Q^{i}_{L}}\widetilde{H}_{2}u_{R}^{j}-Y^{ij}_{d}\overline{Q^{i}_{L}}H_{1}d_{R}^{j}-Y^{ij}_{e}\overline{L^{i}_{L}}H_{1}e_{R}^{j}-Y^{ij}_{n}\overline{L^{i}_{L}}\widetilde{H}_{2}\nu_{R}^{j}+h.c.,$ (23)
while in the Type-Y 2HDM the Yukawa interaction is given by
${\cal L}_{\rm Yukawa}^{Y}=-Y^{ij}_{u}\overline{Q^{i}_{L}}\widetilde{H}_{2}u_{R}^{j}-Y^{ij}_{d}\overline{Q^{i}_{L}}H_{1}d_{R}^{j}-Y^{ij}_{e}\overline{L^{i}_{L}}H_{2}e_{R}^{j}-Y^{ij}_{n}\overline{L^{i}_{L}}\widetilde{H}_{1}\nu_{R}^{j}+h.c.,$ (24)
where, $\widetilde{H}_{1,2}=i\sigma_{2}H_{1,2}^{*}$ and i,j=1, 2, 3i,j=1,\,2,\,3.
In both models, the gauge anomalies involving $U(1)_{H}$ currents are not cancelled with SM fermions only, and additional new chiral fermions have to be introduced in order to fulfill the anomaly cancellation conditions. We could find a lot of setups without gauge anomaly Ko:2012hd . The Type-II 2HDM, for instance, can be realized by the model inspired by the $E_{6}$ Grand Unified Theory Ko:2015fxa . The anomaly-free conditions in the Type-Y 2HDM can be satisfied, by adding extra quarks and leptons in Table 4 when the $U(1)_{H}$ charge assignments to the SM fermions are as in Table 3. Note that Yukawa couplings between $\Phi$ and extra fermions are allowed by the assumed gauge symmetries:
${\cal L}^{Y}_{{\rm extra}}=-y^{i}_{D}\overline{u^{\prime\,i}_{L}}\Phi^{\dagger}u^{\prime\,i}_{R}-y^{i}_{L}\overline{e^{\prime\,i}_{L}}\Phi e^{\prime\,i}+h.c..$ (25)
$u^{\prime\,i}_{L}$ $u^{\prime\,i}_{R}$ $e^{\prime\,i}_{L}$ $e^{\prime\,i}_{R}$
$SU(3)_{c}$ 3 3 1 1
$SU(2)$ 1 1 1 1
$U(1)_{Y}$ $2/3$ $2/3$ 1-1 1-1
$U(1)_{H}$ 11 0 1-1 0
Table 4: The charge assignment to the extra fermions in the Type-Y model.
In this paper, we do not pay attention to the extra fermions too much because they strongly depend on the model construction which is not unique. Our motivation is to study the effects of the extra fermions to the scalar boson decays such as $\tilde{h}\rightarrow\gamma\gamma$ through the loops. Thus, the extra fermion contributions to the loop diagrams must be considered collectively, that is the sum of all the fermions in the model, which is quite cumbersome. Instead, we assume a vectorlike charged lepton and quark as a representative of collection of the extra fermions in the loop, denoting the Yukawa couplings of extra quarks and leptons as $y_{D}$ and $y_{L}$, respectively. In our phenomenological analysis, those parameters are fixed at $y_{D}=y_{L}=1$.
3 Constraints
In this section, we consider various theoretical and phenomenological constraints on our models.
3.1 Theoretical constraints
First, we consider the constraints from the vacuum stability conditions for nonzero VEVs. The requirement for the scalar potential to be stable and bounded from below constrains the dimensionless couplings in the scalar potential. Following the approach in Ref. Arhrib:2011uy , we find the constraints on the dimensionless couplings, which are presented in appendix A.
The models also have constraints from perturbative unitarity bounds. Following the method to find the constraints from perturbative unitarity Horejsi:2005da ; Muhlleitner:2016mzt , we find that the constraints are given by
$|b_{\pm}|,|c_{\pm}|,|f_{\pm}|,|g_{\pm}|,|f_{s,s_{1},s_{2}}|,\frac{1}{2}|a_{1,2,3}|\leq 8\pi,$ (26)
where
$b_{\pm}$ $=$ $2\lambda_{1,2},\quad c_{\pm}=\lambda_{1}+\lambda_{2}\pm\sqrt{(\lambda_{1}-\lambda_{2})^{2}+\lambda_{4}^{2}},$
$f_{\pm}$ $=$ $\lambda_{3}+\lambda_{4}\pm\lambda_{4},\quad g_{\pm}=\lambda\pm\lambda_{4},$
$f_{s}$ $=$ $2\lambda_{\Phi},\quad f_{s_{1}}=\tilde{\lambda}_{1},\quad f_{s_{2}}=\tilde{\lambda}_{2},$ (27)
and $a_{1,2,3}$ are roots of the cubic equation for $x$:
$0$ $=$ $x^{3}-4(3\lambda_{1}+3\lambda_{2}+2\lambda_{\Phi})x^{2}$
- $4(2\tilde{\lambda}_{1}^{2}+2\tilde{\lambda}_{2}^{2}+36\lambda_{1}\lambda_{2}-4\lambda_{3}^{2}-4\lambda_{3}\lambda_{4}-\lambda_{4}^{2}+24\lambda_{1}\lambda_{\Phi}+24\lambda_{2}\lambda_{\Phi})x$
$+$ $96\tilde{\lambda}_{1}^{2}\lambda_{2}+96\tilde{\lambda}_{2}^{2}\lambda_{1}-64\tilde{\lambda}_{1}\tilde{\lambda}_{2}\lambda_{3}-1152\lambda_{1}\lambda_{2}\lambda\Phi+128\lambda_{3}^{2}\lambda_{\Phi}+128\lambda_{3}\lambda_{4}\lambda_{\Phi}+32\lambda_{4}^{2}\lambda_{\Phi}.$
3.2 Electroweak precision observables
In order to study the physical effects on the electroweak precision observables (EWPOs) in this model, full calculations of the relevant amplitudes at the one-loop level are required, which are quite involved. Instead, we calculate the $\Delta T$ parameter approximately. Due to the $ZZ^{\prime}$ mixing, the mass of the $Z$ boson is shifted
$m_{Z}^{2}=\frac{m_{W}^{2}}{c_{W}^{2}}{\cos^{2}\theta}-m_{Z^{\prime}}^{2}\frac{\sin^{2}{\theta}}{\cos^{2}{\theta}},$ (28)
at tree level. We note that the $W^{\pm}$ mass is not shifted by the $U(1)_{H}$ gauge symmetry. Since the mixing angle is expected to be small, this mass shift changes the $\rho$ parameter
$\Delta\rho_{\theta}=1-\frac{1}{\rho}\approx-\sin^{2}\theta\left(1-\frac{c_{W}^{2}m_{Z^{\prime}}^{2}}{m_{W}^{2}}\right)$ (29)
up to the leading order of $s_{\theta}$. This is converted to the $T$ parameter as $T_{\theta}=\Delta\rho_{\theta}/\alpha(m_{Z})$.
The new scalar bosons also contribute to the shift of the $\rho$ parameter at loop levels. At one-loop level, we find that the approximate formula of the contribution is Grimus:2007if
$T_{s}$ $=$ $\frac{1}{16\pi m_{W}^{2}s_{W}^{2}}\left[g_{W^{\pm}H^{\mp}A}^{2}F(m_{H^{\pm}}^{2},m_{A}^{2})+\sum_{i}g_{W^{\pm}H^{\mp}S_{i}}^{2}F(m_{H^{\pm}}^{2},m_{S_{i}}^{2})\right$. (30)
$\qquad\qquad\quad-\sum_{i}g_{ZAS_{i}}^{2}F(m_{A}^{2},m_{S_{i}}^{2})+3\sum_{i}g_{ZZS_{i}}^{2}F(m_{Z}^{2},m_{S_{i}}^{2})$
$\qquad\qquad\quad-3\sum_{i}g_{WWS_{i}}^{2}F(m_{W}^{2},m_{S_{i}}^{2})+3\sum_{i}g_{ZZ^{\prime}S_{i}}^{2}F(m_{Z^{\prime}}^{2},m_{S_{i}}^{2})$
$\qquad\qquad\quad-3g_{W^{\pm}H^{\mp}Z^{\prime}}^{2}F(m_{Z^{\prime}}^{2},m_{H^{\pm}}^{2})-3g_{W^{\pm}H^{\mp}Z}^{2}F(m_{Z}^{2},m_{H^{\pm}}^{2})$
$\left.\qquad\qquad\quad-3F(m_{Z}^{2},m_{h}^{2})+3F(m_{W}^{2},m_{h}^{2})\right],$
where the loop function $F(x,y)$ is defined by
$F(x,y)=\frac{x+y}{2}-\frac{xy}{x-y}\ln\frac{x}{y}$ (31)
for $x\neq y$ while $F(x,y)=0$ for $x=y$. The effective couplings, $g_{W^{\pm}H^{\mp}A}$ and so on, are defined in Appendix B. This loop contribution to $T$ in the case of 2HDMs with $Z_{2}$ symmetry vanishes. However, in our model, the EW symmetry breaking makes the $Z^{\prime}$ boson massive by absorbing one of degrees of freedom in the scalar fields. Therefore, the presence of the $U(1)_{H}$ gauge symmetry implies that the contribution from the $ZZ^{\prime}$ mixing must be taken into consideration together with the scalar loop corrections. Finally, the sum of two contributions leads to
$T=T_{\theta}+T_{s},$ (32)
where the measured value is ParticleDataGroup:2024cfk
$T^{\rm exp}=0.04\pm 0.03.$ (33)
We constrain the $T$ parameter within $2\sigma$.
3.3 Flavor physics
The charged Higgs boson mass $m_{H^{\pm}}$ and $\tan\beta$ are strongly constrained by flavor physics, in particular, by the $b\to s\gamma$ decay. The constraints on the 2HDMs from flavor physics including the experimental results at the LHC were analyzed in Ref. Arbey:2017gmh ; Haller:2018nnx . In the Type-II 2HDM, the lower bound of the charged Higgs mass is about 600 GeV, and this is almost independent of $\tan\beta$. The study of $B\to X_{s}\gamma$ in Ref. Misiak:2017bgg also has led the same lower bound. In Ref. Misiak:2020vlo , the authors improve the SM prediction of $B\to X_{s}\gamma$: $Br(B\to X_{s}\gamma)=(3.40\pm 0.17)\times 10^{-4}$ Misiak:2020vlo . They also comment that the bound of $m_{H^{\pm}}$ becomes very severe: $m_{H^{\pm}}\gtrsim 800$ GeV Misiak:2020vlo . [^1] See also Ref. Li:2024kpd . One of the reasons why the bound becomes so strong is that the improved SM prediction is much higher than the experimental result, $Br(B\to X_{s}\gamma)=(3.32\pm 0.15)\times 10^{-4}$ HFLAV:2016hnz . In the Type-II 2HDM, the charged Higgs contribution to $B\to X_{s}\gamma$ interferes with the SM contribution in a constructive manner, so that such new physics contribution is strongly limited.
Recently, the Heavy Flavor Averaging Group (HFLAV) announced the new result, $Br(B\to X_{s}\gamma)=(3.49\pm 0.19)\times 10^{-4}$ HeavyFlavorAveragingGroupHFLAV:2024ctg , that is a little larger than the SM prediction. If the latest HFLAV result is used, the $m_{H^{\pm}}$ bound is expected to be relaxed. We calculate $Br(B\to X_{s}\gamma)$ using SuperIso v4.1 Mahmoudi:2008tp and compare the new HFLAV result. We derive the lower bound of the charged Higgs mass as $m_{H^{\pm}}\gtrsim 500$ GeV, requiring that our prediction is within $2\sigma$ of the new HFLAV result. The more detailed study is beyond our scope, but more accurate analysis of $B\to X_{s}\gamma$ may improve the bound.
Since both bounds are meaningful, we present plots for two cases: $m_{H^{\pm}}>500$ GeV and $m_{H^{\pm}}>800$ GeV. We refer to the former bound as the HFLAV cut and the latter as the tight cut. We note that the bound from $R_{\gamma}=Br_{(s+d)\gamma}/Br_{cl\bar{\nu}}$ is stronger than that from $Br(B\to X_{s}\gamma)$ Li:2024kpd . Therefore, for the tight cut, we apply the constraints on $m_{H^{\pm}}$ and $\tan\beta$ as provided in Ref. Li:2024kpd , where both bounds are simultaneously considered. However, the HFLAV group has not yet reported a value for $R_{\gamma}$. In Ref. Li:2024kpd , the bounds from $Br(B\to X_{s}\gamma)$ and $R_{\gamma}$ differ by about 50 GeV in the lower limit on $m_{H^{\pm}}$. Taking this difference into account, we present plots using the constraint $m_{H^{\pm}}>500$ GeV for the HFLAV cut. In addition, we have tested our models with a more conservative constraint, $m_{H^{\pm}}>550$ GeV, motivated by the bound from $R_{\gamma}$. However, we find that the resulting behavior is not significantly affected by the 50 GeV difference in the $m_{H^{\pm}}$ bound.
Another stringent constraint from flavor physics arises from the $B_{s}\to\mu^{+}\mu^{-}$ decay. In the Type-II 2HDM, this constraint is significant in both the low and high $\tan\beta$ regimes, i.e., $\tan\beta\sim 1$ and $\tan\beta\sim 20$. In contrast, in the Type-Y 2HDM, only the low $\tan\beta$ regime is affected. We apply these constraints to both the HFLAV and tight cuts as provided in Ref. Li:2024kpd . In our model, there are extra quarks that may lead destructive interference with the SM one. The study of flavor physics concerned with the extra quarks would not be simple because the other flavor observables as well as the collider constraints come into play. The detailed study of flavor physics is our future work.
3.4 Experiments for scalar bosons
There have been a lot of searches for new scalar bosons at the LEP, Tevatron, and LHC, and lots of data have been accumulated. However, there is no definite signal for a new scalar boson so far. These search data except for the resonance data around $96$ GeV will strongly constrain our model. We apply these experimental constraints for the additional scalar bosons using the public code HiggsBounds Bechtle:2020pkv which is encoded in HiggsTools Bahl:2022igd .
In our model, the 125 GeV scalar boson will mix with other neutral scalar bosons. Therefore, these mixings change the predictions for the properties of the observed Higgs boson. Then the precision observables for productions and decays of the 125 GeV Higgs boson also constrain our model strongly. We use the public code HiggsSignals Bechtle:2020uwn to test the 125 GeV Higgs boson in our model. In analysis, we utilize HiggsTools Bahl:2022igd which contains the most recent version of HiggsSignals. In order to ensure that the 125 GeV Higgs boson in our model mimics the behavior of the SM Higgs boson, we impose the conditions $\Delta\chi_{125}^{2}<6.18$ and $\Delta\chi^{2}<0$ Chen:2023bqr , where $\Delta\chi_{125}^{2}$ is the difference of $\chi^{2}$ for the measurement of the 125 GeV Higgs boson between our model and the SM, while $\Delta\chi^{2}$ is that of the combined $\chi^{2}$ of the measurements at the 96 GeV excess as well as those of the 125 GeV Higgs Boson.
3.5 Constraints from $t\bar{t}\tau^{+}\tau^{-}$ production at the LHC
The 96 GeV scalar boson couples to both the top quark and tau lepton. Thus, the production of $t\bar{t}\tau^{+}\tau^{-}$ at the LHC can constrain the scalar boson Iguro:2022dok  [^2] See also Ref. Iguro:2022fel .. Based on the ATLAS Run 2 full data ATLAS:2022yrq , the authors in Ref. Iguro:2022dok have proposed the upper bounds of the Yukawa couplings, $\rho^{tt,\tau\tau}_{\tilde{h}}$, normalized by $\sqrt{2}$. Recently, the CMS collaboration also announced the result of the new physics search in the $t\bar{t}\tau^{+}\tau^{-}$ channel CMS:2024ulc , and the upper bound of the cross section is about $0.6$ times lower than those in Ref.Iguro:2022dok . Using the CMS results, we require that the Yukawa couplings should satisfy
$1.07\times\left|\rho_{\tilde{h}}^{tt}\right|^{2}{\rm Br}\left(\widetilde{h}\to\tau\overline{\tau}\right)$ $<0.03$ , (34)
where $\rho_{\tilde{h}}^{tt}=\frac{g_{\tilde{h}}^{t,\textrm{Y}}m_{t}}{v}$ is defined. It is worthwhile to mention that these constraints are based on the analysis of the tree-level calculation for the $t\bar{t}\tau^{+}\tau^{-}$ production Iguro:2022dok . With QCD corrections to the process, the bounds may be slightly modified.
4 Results
In this section, we analyze the resonance signals in both the Type-II 2HDM and the Type-Y 2HDM.
4.1 Parameters
In our model, there are 11 parameters as follows:
$\alpha_{1},\alpha_{2},\alpha_{3},\tan\beta,v_{\Phi},m_{H},m_{A},m_{H^{\pm}},y_{D},y_{L},g_{X},$ (35)
which denote three mixing angles between CP-even neutral Higgs bosons, $v_{2}/v_{1}$, the VEV of $\Phi$, the heavy neutral Higgs boson mass, the pseudoscalar boson mass, the charged Higgs boson mass, the Yukawa couplings of extra quark and lepton, and the gauge coupling of the $U(1)_{H}$ symmetry, respectively.
For the three mixing angles, we allow the entire range, $0\leq\alpha_{1,2,3}\leq 2\pi$. For $\tan\beta$ and $v_{\Phi}$, we use the following ranges
$1\leq\tan\beta\leq 30\quad\textrm{and}\quad 5\times 10^{3}\leq v_{\Phi}/\textrm{GeV}\leq 10^{4}.$ (36)
The ranges of the masses of the extra scalar bosons are taken to be $500\leq m_{H,A,H^{\pm}}/$GeV $\leq 1500$, respectively, considering the bounds from flavor physics. We have checked broader ranges for $\tan\beta$, $v_{\Phi}$, and the scalar boson masses, but we have found that it does not change the overall behavior of signal strengths in allowed regions. We have also limited the angles in the ranges of $0.5\leq\alpha_{1}\leq 1.5$, $1\leq\alpha_{2}\leq 2$, and $1\leq\alpha_{3}\leq 2$ because no allowed regions exist outside these ranges in the first scan with the entire range, or the other allowed regions does not change the overall feature. At first, the Yukawa couplings of the extra quark and lepton are in the ranges of $0\leq y_{D,L}\leq 4\pi$, respectively, but for simplicity we set $y_{D}=y_{L}=1$. We note that essential features from the analysis are not modified even though larger Yukawa couplings are chosen. The $U(1)_{H}$ gauge coupling must not be large because of constraints from the $Z^{\prime}$ boson search at the LHC. We take $g_{X}$ to be in the range of $0.01\leq g_{X}\leq 0.1$. Then, the $Z^{\prime}$ mass, $m_{Z^{\prime}}$, is written in terms of the parameters
$m_{Z^{\prime}}=\frac{g_{H}v}{2},$ (37)
where
$g_{H}=\sqrt{4g_{X}^{2}\left(\sin^{2}\beta+\frac{v_{\Phi}^{2}}{v^{2}}\right)}.$ (38)
We find that the $Z^{\prime}$ mass is in the range of $100\sim 1000$ GeV for allowed parameters while the VEV for the $U(1)_{H}$ symmetry breaking is in the range of $5\sim 10$ TeV.
Figure 1: Signal strengths $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$ in the Type-II models with extra fermion contributions in the loop (left) and without them (right).
Figure 2: Signal strengths $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$ in the Type-Y models with extra fermion contributions in the loop (left) and without them (right).
4.2 Analysis
We perform a scan of 11 parameters in the ranges described in the previous subsection. Then, we apply the theoretical and experimental constraints to the parameter sets and calculate $\mu_{\gamma\gamma},\mu_{\tau\tau}$ and $\mu_{bb}$ for those parameters that pass those constraints.
In Fig. 1, we present the signal strengths $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$ in both Type-II models with extra fermion contributions in the loop (left) and without those contributions (right), respectively. The extra fermion contribution in the loop can especially modify branching ratios of $\tilde{h}\to gg$ or $\gamma\gamma$. Thus, the production or decay rates of $\tilde{h}$ can be affected by the extra fermion loop contributions. The two dashed lines denote $1\sigma$ and $2\sigma$ regions for the combined experimental values in ATLAS and CMS, respectively. The blue and red points represent parameter sets consistent with the constraints discussed in the previous section. The only distinction between these two sets lies in the flavor physics constraints applied. The blue points satisfy the HFLAV cut for the $B\to X_{s}\gamma$ decays, which imposes a lower bound of $m_{H^{\pm}}\gtrsim 500$ GeV. Since the $B_{s}\to\mu^{+}\mu^{-}$ constraint is more stringent than those from other observables in both low and high $\tan\beta$ regimes Li:2024kpd , we impose the bound provided in Ref. Li:2024kpd . The red points satisfy the tight cut, which is based on the constraints in Ref. Li:2024kpd . We note that the lower limit of $m_{H^{\pm}}$ in this cut is about 800 GeV while the $B_{s}\to\mu^{+}\mu^{-}$ constraint is the same as those in the HFLAV cut. Since the tight cut is more stringent than the HFLAV cut, all the red points overlap the blue points. The yellow points represent parameter sets that satisfy all constraints except for those from HiggsSignals, while adhering to the HFLAV cut for the flavor physics constraints. We note that the constraints from $t\bar{t}\tau^{+}\tau^{-}$ production at the LHC are applied to both red and blue points, while they are not to the yellow points.
We find that the expectation in our Type-II models is inconsistent with the experimental results within $1\sigma$, while some regions are consistent within $2\sigma$. In particular, it is difficult to enhance the $\mu_{\tau\tau}$ because of effective Yukawa couplings of the 96 GeV scalar boson to the top quarks and tau lepton in the Type-II models:
$g_{\tilde{h}}^{t,\textrm{II}}=\frac{\sin\alpha_{1}\cos\alpha_{2}}{\sin\beta},~~~g_{\tilde{h}}^{\tau,\textrm{II}}=\frac{\cos\alpha_{1}\cos\alpha_{2}}{\cos\beta},$ (39)
which are normalized to the SM-like Yukawa couplings, respectively. For the $gg\to\tilde{h}\to\tau\tau$ production, both Yukawa couplings are relevant. Then, $g_{\tilde{h}}^{t,\textrm{II}}$ is proportional to $\sin\alpha_{1}$, but $g_{\tilde{h}}^{\tau,\textrm{II}}$ is to $\cos\alpha_{1}$. Therefore, it would be difficult to enhance both Yukawa couplings for the same $\alpha_{1}$. It turns out that the branching ratio for $\tilde{h}\to\tau\tau$ is at most $0.1$ in the allowed parameter space, which results in the small signal strength $\mu_{\tau\tau}$.
As we see in Fig. 1, the region allowed by the HFLAV cut is broader than that consistent with the tight cut, in the Type-II models. However, the total signal strengths do not differ significantly in the two cases. We find that the constraint from HiggsSignals is very stringent, excluding substantial portion of the parameter space. Without these constraints, the model prediction could deviate by up to approximately $1.3\sigma$ from the experimental central values. This situation is similar in both models, with and without the extra fermion contributions in the loop. Despite this, it is apparent that the signal strengths are distinctive in both models. This implies that a parameter set consistent within $2\sigma$ with the extra fermion contributions might be excluded if the extra fermions are ignored in the loop, and vice versa. Thus, the extra fermions may play a role in the search for the 96 GeV boson.
In Fig. 2, we present the signal strengths $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$ in both Type-Y models with extra fermion contributions in the loop (left) and without those contributions (right), respectively. In these models, the effective Yukawa couplings are given by
$g_{\tilde{h}}^{t,\textrm{Y}}=g_{\tilde{h}}^{\tau,\textrm{Y}}=g_{\tilde{h}}^{t,\textrm{II}}.$ (40)
Since the two effective Yukawa couplings are identical, both production cross sections for $gg\to\tilde{h}\to\gamma\gamma$ and $gg\to\tilde{h}\to\tau\tau$ tend to increase simultaneously as $\sin\alpha_{1}$ increases. This behavior is indeed observed for both the blue and red points, which adhere to the HFLAV and tight cuts, respectively. Moreover, compared to the Type-II models, the $\mu_{\tau\tau}$ signal strength can be enhanced and the deviation from the experimental central values can fall within the $1\sigma$ range under the HFLAV cut. Even under the tight cut, the deviation can reach up to approximately $1.2\sigma$ from the central values. As shown in Fig. 2, the effects of extra fermions in the loop are very restrictive, but one can observe a small difference between two models with and without extra fermions in the loop. We find that some points which are excluded the model without extra fermions could be consistent within $2\sigma$ by including the extra fermions in the loop, similar to the Type-II models.
Figure 3: Signal strengths $\mu_{bb}$ and $\mu_{\tau\tau}$ in the Type-II model with extra fermions in the loop (left) and without them (right).
Figure 4: Signal strengths $\mu_{bb}$ and $\mu_{\tau\tau}$ in the Type-Y model with extra fermions in the loop (left) and without them (right).
In Figs. 3 and 4, we present the signal strengths $\mu_{bb}$ and $\mu_{\tau\tau}$ in the Type-II and Type-Y model 2HDMs, respectively. The extra fermion contributions are included in the left figures, while they are not included in the right figures. The effective couplings of $\tilde{h}$ to the bottom quark are
$g_{\tilde{h}}^{b,\textrm{II}}=g_{\tilde{h}}^{b,\textrm{Y}}=g_{\tilde{h}}^{\tau,\textrm{II}}.$ (41)
In the Type-II 2HDM, the effective Yukawa coupling, $g_{\tilde{h}}^{b,\textrm{II}}$, of $\tilde{h}$ to the bottom quark is the same as that to the tau lepton. Thus, $\mu_{bb}$ tends to be proportional to $\mu_{\tau\tau}$ as shown in the left figure of Fig. 3. It is obvious that $\mu_{bb}$ in the Type-II models is preferred to be similarly small like $\mu_{\tau\tau}$ in the allowed parameter space. On the other hand, in the Type-Y models, the effective Yukawa coupling, $g_{\tilde{h}}^{b,\textrm{Y}}$, of $\tilde{h}$ to the bottom quark is different from that to the tau quark. As shown in Fig. 4, a wider region is allowed in the Type-Y models.
In our models, the extra fermions are required to ensure anomaly cancellation. These fermions can interact with both gauge bosons and scalar bosons. As a result, the branching ratios of scalar bosons into gauge bosons can be affected by extra fermions involving loop diagrams. However, as shown in the figures in this section, the overall features of the models remain unchanged by the inclusion of extra fermion contributions. Therefore, the impact of the extra fermions appears to be insignificant for the 96 GeV resonances. This is because the parameters in the scalar potential are tightly constrained by experimental data, such as flavor physics observables and precision measurements of the SM Higgs boson. The extra fermions play only a minor role in both types of experiments. However, we note that certain detailed features of the models may differ depending on whether the extra fermions are included. The signal strengths, $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$ depend on the branching ratios of $\tilde{h}$ decays into $gg$ and/or $\gamma\gamma$. Therefore, the extra fermion contributions to these branching ratios may become significant if a given parameter point lies near the boundary of the $2\sigma$ allowed region.
4.3 Benchmark points
[…]
There is no big difference between in the Type-II 2HDM and in the Type-Y 2HDM, when charged Higgs is heavy, i.e., $m_{H^{\pm}}>800$ GeV. We note that the parameter values of point 1 in Table 5 are the same as them of point 3. We can see the difference of the mixing dependence on the $\tau$ couplings in the signal strengths. The mixing angles, $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{3}$, in all cases, correspond to ${\cal O}(1)$ $|\sin\alpha_{1,2,3}|$. Then, $\tilde{h}$, whose mass is 96 GeV, is dominated by $h_{\Phi}$, while the SM-like Higgs boson, $h$, dominantly consists of $h_{1}$ and $h_{2}$. In all points in Table 5, $\tan\beta$ is small, and $|\cos\alpha_{1,2}|$ are larger than $0.1$ to enhance $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$. As shown in Table 5, $g_{X}$ is very small to evade too large deviation of the $\rho$ parameter, so, the mixing parameters are relevant to the signal strengths.
point 1 point 2 point 3 point 4
Model Type-II Type-II Type-Y Type-Y
$\alpha_{1}$ 1.01 0.97 1.01 1.36
$\alpha_{2}$ 1.88 1.25 1.88 1.90
$\alpha_{3}$ 1.40 1.48 1.40 1.30
$\tan\beta$ 1.20 1.21 1.20 1.98
$v_{\Phi}$ (GeV) 7459.47 8646.47 7459.47 9899.79
$m_{H}$ (GeV) 900.36 749.49 900.36 652.26
$m_{A}$ (GeV) 713.55 630.49 713.55 549.93
$m_{H^{\pm}}$ (GeV) 877.49 692.21 877.49 605.38
$y_{D}$ 1.0 1.0 1.0 1.0
$y_{L}$ 1.0 1.0 1.0 1.0
$g_{X}$ $3.83\times 10^{-2}$ $7.29\times 10^{-2}$ $3.83\times 10^{-2}$ $1.61\times 10^{-2}$
$\mu_{\gamma\gamma}$ 0.13 0.15 0.13 0.27
$\mu_{\tau\tau}$ $9.07\times 10^{-2}$ 0.13 0.15 0.35
$\mu_{bb}$ $8.59\times 10^{-2}$ $9.25\times 10^{-2}$ $8.08\times 10^{-2}$ $5.23\times 10^{-2}$
Table 5: Benchmark points in the Type-II and Type-Y 2HDMs with $U(1)_{H}$ gauge symmetry.
5 Summary
We studied the di-photon and di-$\tau$ channels at the LHC as well as the $b\bar{b}$ channel at LEP, in the Type-II and Type-Y 2HDMs with gauged $U(1)_{H}$ symmetry. We can simply add one extra Higgs doublet to the SM. Such a simple extension, however, causes FCNCs at the tree level, if two Higgs doublets are not distinguished. Moreover, extra scalars predicted by the two Higgs doublets usually reside around the EW scale, so that the extended SMs are easily excluded. In the conventional 2HDMs (e.g., the Type-II 2HDM and the Type-Y 2HDM), softly broken $Z_{2}$ symmetry is imposed on fields and the strong constraints are evaded. In our model, we see that softly broken $Z_{2}$ symmetry is originated from the spontaneous symmetry breaking of gauged $U(1)_{H}$ symmetry. In this sense, our model can be interpreted as the underlying theory of the conventional 2HDMs. Moreover, the 2HDM with a singlet scalar is also recently studied, motivated by the excesses around 96 GeV. Our 2HDM with gauged $U(1)_{H}$ symmetry also predicts a singlet scalar, so it is worthwhile to study the signals related to the resonance around 96 GeV in our models. In our model, the $U(1)_{H}$ gauge symmetry distinguishes one Higgs doublet from the other, and naturally suppresses tree-level FCNCs. One extra scalar charged under $U(1)_{H}$ is also introduced, so that there are three CP-even neutral scalars at low energy. In our study, the mass of lightest neutral scalar, denoted by $\tilde{h}$, was fixed at 96 GeV that corresponds to the mass region where the mild excesses are reported by the ATLAS and CMS collaborations. In our models, extra fermions as well as extra scalars contribute to the di-photon and di-$\tau$ channels. The extra gauge boson deviates $\rho$ parameter from $1$. We surveyed our predictions of the signal strengths, $\mu_{bb}$, $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$, scanning parameters. The $Z^{\prime}$ couplings are strongly constrained by the $\rho$ parameter and the $Z^{\prime}$ searches, while the contribution of extra fermions affects the signal strengths in some parameter sets. We saw that the dominant contributions to $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$ are from the mixing among three neutral scalars, not from the extra new fermions. As discussed in Sect. 4, the lightest neutral scalar is mainly composed of the scalar from $\Phi$. This result is consistent with the previous work Biekotter:2023oen . Our work reveals the underlying theory of the 2HDM with a complex scalar field Biekotter:2023oen , and explicitly shows how large contributions of the extra fields predicted by extending the symmetry, that distinguishes two Higgs fields, to gauged $U(1)_{H}$ symmetry can be. The predictions of the signal strengths, in fact, deviate from them of the 2HDMs without extra fermions. Further study would be required to reveal which setup is preferred by the excesses around 96 GeV. The improved SM prediction of $B\to X_{s}\gamma$ may exclude the parameter region with $m_{H^{\pm}}<800$ GeV Misiak:2020vlo , but the new HFLAV result HeavyFlavorAveragingGroupHFLAV:2024ctg would relax the bound drastically. We plot our predictions in both heavy and light charged Higgs cases. There is no big difference between in the Type-II 2HDM and in the Type-Y 2HDM, when charged Higgs is heavy. If the charged Higgs is light, the signal strength, $\mu_{\tau\tau}$, tends to be larger in the Type-Y than that in the Type-II. In our model, there are extra quarks which can contribute to $B\to X_{s}\gamma$. The study of flavor physics involving extra quarks, however, is not simple, since other flavor observables are usually correlated to $B\to X_{s}\gamma$. This is also our future work. Finally, let us comment on how to distinguish our models with the others in the previous works. As we see in Table 5, $Z^{\prime}$ mass, that is ${\cal O}(g_{X}v_{\Phi})$, is below 1 TeV in our model. There are several possibilities of $U(1)_{H}$ charge assignments and matter fields as discussed in Ref. Ko:2012hd , but in any case, the resonance search below 1 TeV play an important role in testing our models. As studied in Ref Ko:2015fxa , there is a good DM candidate in the Type-II 2HDM with gauged $U(1)_{H}$ symmetry. Our analysis in this paper does not include the DM, but this prediction is also important to test and distinguish our model with the others.
Acknowledgements.
This work is supported in part by National Research Foundation of Korea (NRF) Grant No. NRF-2018R1A2A3075605 and No. RS-2023-00270569 (S. B.), KIAS Individual Grants under Grant No. PG021403 (P. K.), Grant-in-Aid for Scientific research from the MEXT, Japan, No. 24K07031 (Y. O.), Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Science, ICT, and Future Planning under the Grant No. NRF-2021R1A2C2011003m and RS-2023-00237615 (C. Y.).
Appendix A Bounded-from-below condition
We present constraints from the condition for scalar potential to be bounded from below. For the derivation of the constraints, we follow the method derived in Ref. Horejsi:2005da ; Muhlleitner:2016mzt . The quartic couplings are constrained by
$\lambda_{1,2,\Phi}>0,\quad\lambda_{3}+(\lambda_{4}\pm\lambda_{4})/2+2\sqrt{\lambda_{1}\lambda_{2}}>0,\quad\tilde{\lambda}_{1,2}+2\sqrt{\lambda_{1,2}\lambda_{\Phi}}>0$
together with the following four conditions as follows
$[(\lambda_{3}>0\cap\lambda_{34}>0)\cup(E_{1})]\cap[(\tilde{\lambda}_{1}>0)\cup(4\lambda_{1}\lambda_{\Phi}-\tilde{\lambda}_{1}^{2}>0)]$ (43)
$\cap$ $[(\tilde{\lambda}_{2}>0)\cup(4\lambda_{2}\lambda_{\Phi}-\tilde{\lambda}_{2}^{2}>0)]\cap[(\tilde{\lambda}_{1}>0\cap\tilde{\lambda}_{2}>)\cup(E_{2})],$
where
$E_{1}$ $=$ $[\lambda_{12}^{2}-\lambda_{3}^{2}>0]\cap[\lambda_{12}^{2}-\lambda_{34}^{2}>0]\cap[\lambda_{12}^{2}-\lambda_{3}\lambda_{34}+\sqrt{(\lambda_{12}^{2}-\lambda_{3}^{2})(\lambda^{2}-\lambda_{34}^{2})}>0],$ (44)
$E_{2}$ $=$ $[\lambda_{1\Phi}^{2}-\tilde{\lambda}_{1}^{2}>0]\cap[\lambda_{2\Phi}^{2}-\tilde{\lambda}_{2}^{2}>0]\cap[2(\lambda_{3}+D)\lambda_{\Phi}-\tilde{\lambda}_{1}\tilde{\lambda}_{2}+\sqrt{(\lambda_{1\Phi}^{2}-\tilde{\lambda}_{1}^{2})(\lambda_{2\Phi}^{2}-\tilde{\lambda}_{2}^{2})}>0].$ (45)
Here,
$\lambda_{12}^{2}$ $=$ $4\lambda_{1}\lambda_{2},\quad\lambda_{1\Phi}^{2}=4\lambda_{1}\lambda_{\Phi},\quad\lambda_{2\Phi}^{2}=4\lambda_{2}\lambda_{\Phi},$ (46)
$D$ $=$ $\min(\lambda_{4},0).$ (47)
Appendix B Effective couplings
In this appendix, we present the definition of the effective couplings for the triple vertices of gauge bosons and scalar bosons. The effective couplings are defined by the couplings for the triple vertices normalized by those in the SM
$g_{W^{\pm}H^{\mp}A}$ $=$ $(V_{A})_{33}\cos\beta-(V_{A})_{32}\sin\beta,$ (48)
$g_{W^{\pm}H^{\mp}S_{i}}$ $=$ $(R)_{i2}\cos\beta-(R)_{i1}\sin\beta,$ (49)
$g_{ZAS_{i}}$ $=$ $\cos\theta\left[(R)_{i1}(V_{A})_{32}+(R)_{i2}(V_{A})_{33}\right]$ (50)
- $\frac{2g_{X}}{g_{2}}\sin\theta\left[(R)_{i2}(V_{A})_{33}-(R)_{i3}(V_{A})_{31}\right],$
$g_{ZZS_{i}}$ $=$ $\cos^{2}\theta\left[(R)_{i1}\cos\beta+(R)_{i2}\sin\beta\right]-\frac{4c_{W}^{2}g_{X}^{2}}{g_{2}^{2}}\sin\theta(R)_{i2}\sin\beta$ (51)
$+$ $\frac{4c_{W}^{2}g_{X}^{2}}{g_{2}^{2}}\sin^{2}\theta\left[(R)_{i2}\sin\beta+(R)_{i3}\frac{v_{\Phi}}{v}\right],$
$g_{ZZ^{\prime}S_{i}}$ $=$ $\frac{2c_{W}g_{X}}{g_{2}}(R)_{i2}\sin\beta(\cos^{2}\theta-\sin^{2}\theta)$ (52)
$+$ $\frac{\sin 2\theta}{2}\left[(R)_{i1}\cos\beta+(R)_{i2}\sin\beta+\frac{4c_{W}^{2}g_{X}^{2}}{g_{2}^{2}}\left\{(R)_{i2}\sin^{2}\beta+\frac{v_{\Phi}}{v}(R)_{i3}\right\}\right],$
$g_{WWS_{i}}$ $=$ ${(R)_{i1}\cos\beta+(R)_{i2}\sin\beta},$ (53)
$g_{W^{\pm}H^{\mp}Z}$ $=$ $\frac{g_{X}}{g_{2}}\sin\theta\sin 2\beta,$ (54)
$g_{W^{\pm}H^{\mp}Z^{\prime}}$ $=$ $-\frac{g_{X}}{g_{2}}\cos\theta\sin 2\beta,$ (55)
where $S_{i}\equiv(\tilde{h},h,H)$.